# RankByGene: Gene-Guided Histopathology Representation Learning Through Cross-Modal Ranking Consistency

### v1

@@ Abstract
Spatial transcriptomics (ST) provides essential spatial context by mapping gene expression within tissue, enabling detailed study of cellular heterogeneity and tissue organization. However, aligning ST data with histology images poses challenges due to inherent spatial distortions and modality-specific variations. Existing methods largely rely on direct alignment, which often fails to capture complex cross-modal relationships. To address these limitations, we propose a novel framework that aligns gene and image features using a ranking-based alignment loss, preserving relative similarity across modalities and enabling robust multi-scale alignment. To further enhance the alignment’s stability, we employ self-supervised knowledge distillation with a teacher-student network architecture, effectively mitigating disruptions from high dimensionality, sparsity, and noise in gene expression data. Extensive experiments on gene expression prediction and survival analysis demonstrate our framework’s effectiveness, showing improved alignment and predictive performance over existing methods and establishing a robust tool for gene-guided image representation learning in digital pathology.
@@ 1 Introduction
Digital pathology has advanced significantly in recent years, thanks to the availability of large number of digitized slides and rapid development of deep learning. Powerful learning methods have been proposed for image-based prediction at different scales, including individual cells [22, 5, 25], larger structures like glands [29, 69, 75], regions of interest [33, 20], and the whole slide [28, 34, 42, 32, 13]. However, these methods, despite their promise, are fundamentally limited by the amount of information one can extract from a digital slide. Most clinical outcomes are given at a whole slide-level. Due to the heterogeneity of the tissue microenvironment, it is very difficult for a machine learning algorithm to directly connect localized cell morphology and spatial context with clinical outcomes. Connecting cell morphology/context with their actual functionality is the missing link and can be the key to advancing image-based diagnosis and prognosis.
Transcriptomics [49, 72] reflects the genome-wide quantification of gene expression within a tissue, and these approaches provide a snapshot of gene activity within a variably heterogeneous collection of cells. Single-cell RNA sequencing [63, 62] attempts to address this cellular heterogeneity by profiling gene expression within individual level, enabling more precise cell typing and characterization. Despite its high resolution, this approach disrupts tissue architecture and loses the critical spatial context that defines an organized tissue [58].
Spatial transcriptomics (ST) bridges the gap by enabling gene expression profiling within tissue structures, preserving the spatial context of cellular activity [56]. It measures gene expression at specific locations, allowing for the mapping of cellular states and molecular interactions in their native micro-environments [61, 45]. This spatial dimension is crucial for studying the interactions between cells and their microenvironment, which are central to understanding disease dynamics and therapeutic responses [48, 19]. For example, ST can uncover molecular signatures unique to specific tumor regions, such as the core, invasive margins, and immune-infiltrated zones, enabling a nuanced view of tumor heterogeneity [31].
Figure 1: t-SNE visualization [66] of image representation from different methods. (a) Natural-Image-SSL. (b) Histopath-Image-SSL. (c) Gene-Image CL (d) Gene-Image CL with ranking loss. (e) RankByGene. We first use K-means clustering to assign a class label to each spot based on gene expression values. Then, we apply t-SNE to visualize the image features corresponding to each spot and color them according to their assigned class labels. A higher v-score [57] indicates that the image representation more precisely captures the distribution of gene expression values. Adding ranking loss (d) enhances the separability between image features of different classes. Our method (e), with both ranking loss and distillation shows the highest separability.
@@
ST data provides new opportunities to enhance existing image-based prediction methods with novel biological information. Most ST data have both localized gene expressions and their corresponding histology image patches. One can exploit the two complementary modalities at a fine scale, unveiling the subtle visual cues from cell morphology and context that capture cell functionality. In this paper, we study the problem of learning gene-guided image representations through ST data. In particular, we use the gene expression to better align image features, so that these image features encode cell-functionality-specific visual information. As we will demonstrate in experiments, these gene-guided image features will improve image-based prediction of clinical outcomes. The learned features can be applied broadly to any image-only cohort, benefiting the digital pathology community.
@@
To address these challenges, methods have been proposed to map both image patches and gene expressions into a shared latent feature space. Self-supervised contrastive learning losses have been adopted to ensure the image features and gene features from the same location (called a spot) are mapped together [30]. The standard contrastive loss, however, overlooks the potential similarity between different spots. Xie et al. [73] employ a smoothed variant of the CLIP loss [55] to encourage spots with similar image patches or similar gene expressions to be mapped closer in the feature space. These methods, however, do not fully address the aforementioned challenges and lack an effective global strategy to align multi-modal features.
@@
In this paper, we propose a novel approach to align image and gene features across multiple scales, both locally and globally. The alignment method is designed to be robust to modality-specific distortions, addressing the challenges outlined earlier. Specifically, we introduce an innovative ranking-based alignment loss that captures the relative similarity ordering of feature representations across modalities. By enforcing consistency in cross-modal ranking, we ensure that the similarity relationships between gene expression features are reflected in the corresponding image features. This ranking-based approach offers several key advantages. First, it facilitates alignment across both local and global scales, enabling even distant features to interact and align. This global alignment complements the existing local matching of image and gene features of the same spot [73, 46]. Second, by focusing on matching rankings rather than exact similarity values, our method is more robust to distortions, allowing it to tolerate similarity variations especially between features that are far apart from one another.
@@
To further enhance the robustness of the alignment process, we adopt a self-supervised knowledge distillation approach to stabilize the learning of gene-informed image representations. By utilizing a teacher-student network architecture with both weakly and strongly augmented image patches, we ensure that the student network effectively learns the aligned features. This approach helps the student network remain resilient to potential disruptions arising from imperfections in the gene expressions, such as high dimensionality, sparsity, and noisy or missing values.
@@
Figure 1 shows the t-SNE visualization [66] of image features obtained from various methods: natural image-only self-supervised learning (SSL), histopathology image-only SSL, gene-image contrastive learning (CL), gene-image CL with ranking loss, and our method. Compared to other methods, the image features generated by our approach effectively capture the distribution of gene expression values.
@@
In summary, we propose a novel multi-modal feature alignment method that integrates pathology images and ST data, enabling robust gene-informed image representation learning. Our contributions are three-fold:
@@
  • We introduce a new cross-modal ranking consistency mechanism to enhance multi-modal alignment, achieving alignment at both local and global scales while remaining robust to distortions over long-range distances.
  • We apply a knowledge distillation approach to the multi-modal alignment task, further improving the robustness of the alignment process despite imperfections in gene expression features.
  • Extensive experiments across multiple downstream tasks, including gene expression prediction and survival prediction, demonstrate the superior performance of our framework compared to existing self-supervised alignment methods.
@@
Figure 2: Overview of our RankbyGene framework. The framework begins with WSI Tiling, where WSIs are divided into tiles, each paired with a gene profile. In the feature extraction part, weak and strong augmentations of the image tiles are processed through a teacher encoder and student encoder, while a gene encoder extracts features from the gene profile. The feature alignment stage ensures that weakly and strongly augmented image features are aligned through intra-modal distillation loss and the image and gene features are aligned using gene-image contrastive loss. Meanwhile, our proposed cross-modal ranking consistency loss maintains consistent similarity ranking orders across two modalities.
@@ 2 Related Work
Histopathology Image Analysis. Histopathology image analysis is regarded as the gold standard for cancer diagnosis and treatment [41, 50, 9]. Due to the huge resolution of whole slide images (WSIs) and the difficulty of obtaining patch-level labels, WSI analysis is commonly performed in a weakly supervised setting [18], where only slide-level labels are available. Typical WSI analysis tasks include classification [28, 34, 77, 59, 54, 64, 78], which predicts the presence or absence of a tumor or identifies specific tumor subtypes, and survival prediction [13, 76, 81, 79, 60], which estimates patient mortality risk. Many multiple instance learning (MIL) [28, 34, 77, 53, 39, 82] based frameworks have been proposed to model relationships between patches more effectively, enabling more accurate slide-level predictions.
@@ 3 Methods
Our learning pipeline is illustrated in Figure 2. The training data consists of pairs of image patches and gene expression profiles, with each pair corresponding to a specific tissue spot. Image patches are fed into image encoders, which include both a teacher and a student encoder to facilitate knowledge distillation for gene-guided image feature learning. This setup helps stabilize the image feature learning process during alignment with gene features. Meanwhile, gene expression data is processed by a gene encoder.
Both image and gene features are mapped into a shared latent feature space. To align these features, we employ a contrastive learning loss that ensures the image and gene features of the same tissue spot are closely matched. Additionally, we introduce a cross-modal ranking consistency loss to maintain the same relative ranking of similarity across spots, ensuring that for any given spot, its similarity to other spots is consistent in both image and gene feature spaces.
@@ 3.2 Cross-Modal Ranking Consistency Loss
Our ranking loss is illustrated in Figure 3. For a spot of interest, $p$, we want to ensure the similarity between $p$ and two other spots, $q$ and $r$, are ranked consistently in both image feature and gene feature. For simplicity, we denote by $S^{I}_{p,q}=\text{sim}(i_{p},i_{q})$ and $S^{G}_{p,q}=\text{sim}(g_{p},g_{q})$ the similarity between $p$ and $q$’s image features and gene features. We also denote by $S^{I}_{p,r}$ and $S^{G}_{p,r}$ the image and gene feature similarity between $p$ and $r$. Assume $S^{G}_{p,q}>S^{G}_{p,r}$. In other words, $g_{q}$ is closer to $g_{p}$ compared with $g_{r}$. We would like to ensure the image features $i_{q}$ and $i_{r}$ also maintain the same order. Classic ordinal ranking loss [10] will just require the image feature similarity difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than a fixed positive value $\epsilon$. However, this may be undesirable especially for long interactions. To better preserve the relative positions between features, we require the difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than $S^{G}_{p,q}-S^{G}_{p,r}$. Formally, we define the function
@@
Accelerating the Computation. During training, at each iteration, we will apply the loss to all spots within a mini-batch. $N$ becomes the batch size. However, this can still be computationally expensive, as it compares $O(N^{3})$ pairs of spots. To alleviate the burden, we still enumerate through all $p$’s. But for each $p$, instead of enumerating through all $O(N^{2})$ $(q,r)$ pairs, we only sample $O(N)$ random $(q,r)$ pairs. In particular, taking the list of spots except $p$, $L=(1,\ldots,p-1,p+1,\ldots,N)$. We randomly shuffle the sequence, getting a shuffled list, $L^{\prime}$. We go through the list, $L^{\prime}$, each time taking two consecutive spots as $q$ and $r$. For the last spot, we will pair it with the first spot in $L^{\prime}$. This gives us $N-1$ sample $(q,r)$ pairs, and ensures each spot appears in two of the sampled pairs. This reasonably covers similarity rankings of all the spots, but only using $O(N^{2})$ triplets.

### v2

@@ Abstract
Spatial transcriptomics (ST) provides essential spatial context by mapping gene expression within tissue, enabling detailed study of cellular heterogeneity and tissue organization. However, aligning ST data with histology images poses challenges due to inherent spatial distortions and modality-specific variations. Existing methods largely rely on direct alignment, which often fails to capture complex cross-modal relationships. To address these limitations, we propose a novel framework that aligns gene and image features using a ranking-based alignment loss, preserving relative similarity across modalities and enabling robust multi-scale alignment. To further enhance the alignment’s stability, we employ self-supervised knowledge distillation with a teacher-student network architecture, effectively mitigating disruptions from high dimensionality, sparsity, and noise in gene expression data. Extensive experiments on seven public datasets that encompass gene expression prediction, slide-level classification, and survival analysis demonstrate the efficacy of our method, showing improved alignment and predictive performance over existing methods.
@@ 1 Introduction
Digital pathology has advanced significantly in recent years, thanks to the availability of large number of digitized slides and rapid development of deep learning. Powerful learning methods have been proposed for image-based prediction at different scales, including individual cells [23, 5, 27], larger structures like glands [31, 73, 80], regions of interest [35, 21], and the whole slide [30, 36, 44, 34, 13]. However, despite their promise, these methods are inherently constrained by the information that can be extracted from a digital slide. Bridging the gap between cell morphology, contextual information, and actual functionality remains a critical challenge, one that could be key to advancing image-based diagnosis and prognosis.
Figure 1: t-SNE [70] visualization of image features of different spots in a ST slide. We show the feature learned using different methods, including (a) SSL on natural images, (b) SSL on histopathology images, (c) CL on ST data using InfoNCE loss [32], and (d) RankByGene. Learning with gene information (c) clearly outperforms learning with images alone ((a) and (b)). Furthermore, our method (d) achieves even greater improvement, demonstrating better inter-class separability than (c), thanks to the proposed contributions. Each spot is assigned a label via K-means clustering on gene expression values; the corresponding image features are then color-coded by these labels. We also provide quantitative measures using v-score [61], where a higher value indicates better alignment of image features with gene expression.
Transcriptomics [51, 76] quantifies gene expression across tissues, providing a snapshot of cellular heterogeneity. Single-cell RNA sequencing [67, 66] refines this by profiling individual cells, enabling precise cell characterization but disrupting tissue architecture and losing spatial context [62]. Spatial transcriptomics (ST) overcomes this limitation by preserving spatial organization while mapping gene expression [60]. It captures cellular states and molecular interactions within their native environment [65, 47], crucial for understanding disease dynamics and therapeutic responses [50, 20]. For instance, ST reveals distinct molecular signatures in tumor regions, such as the core, invasive margins, and immune-infiltrated zones, offering deeper insights into tumor heterogeneity [33]. However, its high cost remains a significant barrier to widespread clinical adoption.
@@
Despite the high cost, ST data provides new opportunities to enhance existing affordable image-based prediction methods with novel biological information. Most ST data have both localized gene expressions and their corresponding histology image patches. One can exploit the two complementary modalities at a fine scale, unveiling the subtle visual cues from cell morphology and context that capture cell functionality. As shown in recent studies [12, 32, 48, 77], visual information including the density and spatial organization of various cell types (e.g., epithelial, lymphocyte, fibroblast) correlates with gene pathways and expressions. If correctly distilled, these gene-correlated visual cues can be valuable for downstream tasks.
In this paper, we study the problem of learning gene-guided image representations through ST data. In particular, we use the gene expression to better align image features, so that these image features encode cell-functionality-specific visual information. As we will demonstrate in experiments, these gene-guided image features will improve image-based prediction of clinical outcomes. The learned features can be applied broadly to any image-only cohort, benefiting the digital pathology community.
@@
To address these challenges, methods have been proposed to map both image patches and gene expressions into a shared latent feature space. Self-supervised contrastive learning losses have been adopted to ensure the image features and gene features from the same location (called a spot) are mapped together [32]. The standard contrastive loss, however, overlooks the potential similarity between different spots. Xie et al. [77] employ a smoothed variant of the CLIP loss [59] to encourage spots with similar image patches or similar gene expressions to be mapped closer in the feature space. Similarly, Min et al. [48] adopt a similar approach to enhance feature alignment. These methods, however, do not fully address the aforementioned challenges and lack an effective global strategy to align multi-modal features.
@@
RankByGene. We propose RankByGene, a novel approach to align image and gene features across multiple scales, both locally and globally. The alignment method is designed to be robust to modality-specific distortions, addressing the challenges outlined earlier. Specifically, we introduce an innovative ranking-based alignment loss ensuring that similarity relationships between gene expression features are reflected in the corresponding image features. This ranking-based approach offers several key advantages. First, it facilitates alignment across both local and global scales, enabling even distant features to interact and align. This global alignment complements the existing local matching of image and gene features of the same spot [77, 48]. Second, by focusing on matching rankings rather than exact similarity values, our method is more robust to distortions, allowing it to tolerate similarity variations, especially between features that are far apart from one another.
@@
To further enhance the robustness of the alignment, we adopt a self-supervised knowledge distillation approach to stabilize the learning of gene-informed image representations. By utilizing a teacher-student network architecture with both weakly and strongly augmented image patches, we ensure that the student network effectively learns the aligned features. This approach helps the student network remain resilient to potential disruptions arising from imperfections in ST gene expression measurements, such as high dimensionality, sparsity, and noisy or missing values.
@@
The strength of our method is demonstrated in Figure 1, which shows the t-SNE visualization [70] of image features from various methods: natural image-only self-supervised learning (SSL), histopathology image-only SSL, gene-image contrastive learning (CL) using InfoNCE [32], and RankByGene. Gene-image CL clearly obtains image features better aligned with gene expressions. Our approach further boosts the features using the proposed contributions.
@@
In summary, we propose a novel multi-modal feature alignment method that integrates pathology images and ST data, enabling robust gene-informed image representation learning. Our contributions are threefold:
@@
  • We propose a novel cross-modal ranking consistency mechanism that enhances gene-image alignment at both local and global scales, while maintaining robustness to long-range distortions.
  • We introduce a knowledge distillation module to the gene-image alignment framework, further improving the robustness of the alignment against artifacts in ST data gene expression such as sparsity and noisy measurements.
  • Extensive experiments on gene expression prediction, gene-related slide-level classification and survival analysis tasks show that our method outperforms existing ones, underscoring its robustness and efficacy.
@@
Figure 2: Overview of our RankbyGene framework. The framework begins with WSI Tiling, where WSIs are cut into patches, each paired with a gene spot. In feature extraction, weak and strong augmentations of the patches are processed through a teacher and student encoder, while a gene encoder extracts features from the gene profile. The feature alignment stage ensures that weakly and strongly augmented image features are aligned through intra-modal distillation loss and the image and gene features are aligned using gene-image contrastive loss. Meanwhile, our proposed cross-modal ranking consistency loss maintains consistent similarity ranking across two modalities.
@@ 2 Related Work
Histopathology Image Analysis. Histopathology image analysis is regarded as the gold standard for cancer diagnosis and treatment [43, 53, 9]. Due to the huge resolution of whole slide images (WSIs) and the difficulty of obtaining patch-level labels, WSI analysis is commonly performed in a weakly supervised setting [19], where only slide-level labels are available. Typical WSI analysis tasks include classification [30, 36, 82, 63, 58, 68, 83], which predicts the presence or absence of a tumor or identifies specific tumor subtypes, and survival prediction [13, 81, 86, 84, 64], which estimates patient mortality risk. Multiple instance learning (MIL) [30, 36, 82, 57, 41, 87] based frameworks have been proposed to model relationships between patches more effectively, enabling more accurate slide-level predictions.
@@ 3 Methods
Our learning pipeline is illustrated in Figure 2. The training data consists of pairs of image patches and gene expression profiles, with each pair corresponding to a specific tissue spot. Image patches are fed into image encoders that include both a teacher and a student encoder to facilitate knowledge distillation for gene-guided image feature learning. This setup helps stabilize the image feature learning process during alignment with gene features. Meanwhile, gene expression data are processed by a gene encoder.
Both image and gene features are mapped into a shared latent feature space. To align these features, we employ a contrastive learning loss that ensures the image and gene features of the same tissue spot are closely matched. A main contribution is that we introduce a cross-modal ranking consistency loss, ensuring that for any given spot, its similarity to other spots is consistent in both image and gene feature spaces. This ranking loss ensures robust global alignment even for spots whose features are not close to each other.
@@ 3.2 Cross-Modal Ranking Consistency Loss
Our ranking loss is inspired by the classic ordinal ranking loss [10], but is modified to better address challenges in our gene-image alignment task. As illustrated in Figure 3, for a spot of interest, $p$, we want to ensure the similarity between $p$ and two other spots, $q$ and $r$, are ranked consistently in both image feature and gene feature. For simplicity, we denote by $S^{I}_{p,q}=\text{sim}(i_{p},i_{q})$ and $S^{G}_{p,q}=\text{sim}(g_{p},g_{q})$ the similarity between $p$ and $q$’s image features and gene features. We also denote by $S^{I}_{p,r}$ and $S^{G}_{p,r}$ the image and gene feature similarity between $p$ and $r$. Assume $S^{G}_{p,q}>S^{G}_{p,r}$. In other words, $g_{q}$ is closer to $g_{p}$ compared with $g_{r}$. We would like to ensure the image features $i_{q}$ and $i_{r}$ also maintain the same order. Classic ordinal ranking loss [10] will just require the image feature similarity difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than a fixed positive value $\epsilon$. In practice, however, we often notice that this loss does not work effectively; image features are often more tightly packed compared with gene features, thus just maintaining the ranking of image similarities will not making much changes, at least in the early alignment stages.
To further accelerate the alignment, we propose a modified ranking loss that pushes image features further apart if possible. In our new ranking loss, we forcing similarity among images to be bigger than similarity among corresponding genes; it pushes image features apart while preserving the ranking. Formally, we require the difference $S^{I}_{p,q}-S^{I}_{p,r}$ to be bigger than $S^{G}_{p,q}-S^{G}_{p,r}$. We define
@@
Accelerating the Computation. During training, at each iteration, we will apply the loss to all spots within a mini-batch (batch size $N$). However, this can still be computationally expensive, as it compares $O(N^{3})$ pairs of spots. To alleviate the burden, we still enumerate through all $p$’s. But for each $p$, instead of enumerating through all $O(N^{2})$ $(q,r)$ pairs, we only sample $O(N)$ random $(q,r)$ pairs. In particular, taking the list of spots except $p$, $L=(1,\ldots,p-1,p+1,\ldots,N)$. We randomly shuffle the sequence, getting a shuffled list, $L^{\prime}$. We go through the list, $L^{\prime}$, each time taking two consecutive spots as $q$ and $r$. For the last spot, we will pair it with the first spot in $L^{\prime}$. This gives us $N-1$ sample $(q,r)$ pairs, and ensures each spot appears in two of the sampled pairs. This reasonably covers similarity rankings of all the spots, but only using $O(N^{2})$ triplets.
Image Encoder Breast-ST1 Breast-ST2 Lung-ST
(a) Results for gene expression prediction (top 250 genes among all genes)
ResNet-50 [25] 0.494 $\pm$ 0.004 0.401 $\pm$ 0.006 0.069 $\pm$ 0.004 0.465 $\pm$ 0.003 0.384 $\pm$ 0.008 0.069 $\pm$ 0.008 0.709 $\pm$ 0.013 0.695 $\pm$ 0.019 0.026 $\pm$ 0.009
CTransPath [74] 0.489 $\pm$ 0.005 0.388 $\pm$ 0.012 0.088 $\pm$ 0.007 0.459 $\pm$ 0.012 0.352 $\pm$ 0.013 0.132 $\pm$ 0.009 0.690 $\pm$ 0.012 0.635 $\pm$ 0.017 0.039 $\pm$ 0.005
UNI [15] 0.488 $\pm$ 0.008 0.383 $\pm$ 0.010 0.096 $\pm$ 0.008 0.461 $\pm$ 0.015 0.335 $\pm$ 0.017 0.154 $\pm$ 0.014 0.681 $\pm$ 0.011 0.619 $\pm$ 0.019 0.043 $\pm$ 0.007
ST-Net [24] 0.491 $\pm$ 0.004 0.395 $\pm$ 0.013 0.085 $\pm$ 0.009 0.464 $\pm$ 0.012 0.362 $\pm$ 0.009 0.104 $\pm$ 0.011 0.685 $\pm$ 0.015 0.615 $\pm$ 0.014 0.048 $\pm$ 0.006
HisToGene [56] 0.485 $\pm$ 0.006 0.385 $\pm$ 0.017 0.102 $\pm$ 0.011 0.458 $\pm$ 0.015 0.329 $\pm$ 0.015 0.144 $\pm$ 0.007 0.674 $\pm$ 0.016 0.603 $\pm$ 0.015 0.055 $\pm$ 0.004
HEST-FT [32] 0.482 $\pm$ 0.011 0.379 $\pm$ 0.015 0.125 $\pm$ 0.013 0.452 $\pm$ 0.013 0.324 $\pm$ 0.016 0.207 $\pm$ 0.005 0.661 $\pm$ 0.011 0.594 $\pm$ 0.018 0.078 $\pm$ 0.004
BLEEP [77] 0.477 $\pm$ 0.015 0.378 $\pm$ 0.017 0.138 $\pm$ 0.009 0.454 $\pm$ 0.016 0.326 $\pm$ 0.011 0.212 $\pm$ 0.006 0.669 $\pm$ 0.015 0.589 $\pm$ 0.012 0.081 $\pm$ 0.001
RankByGene 0.472 $\pm$ 0.010 0.371 $\pm$ 0.012 0.185 $\pm$ 0.005 0.447 $\pm$ 0.008 0.318 $\pm$ 0.010 0.231 $\pm$ 0.012 0.648 $\pm$ 0.013 0.570 $\pm$ 0.011 0.097 $\pm$ 0.006
(b) Results for gene expression prediction (top 250 genes in cancer-specific gene list)
ResNet-50 [25] 0.381 $\pm$ 0.005 0.291 $\pm$ 0.008 0.089 $\pm$ 0.008 0.432 $\pm$ 0.006 0.336 $\pm$ 0.005 0.112 $\pm$ 0.011 0.317 $\pm$ 0.006 0.215 $\pm$ 0.007 0.041 $\pm$ 0.012
CTransPath [74] 0.375 $\pm$ 0.007 0.287 $\pm$ 0.005 0.098 $\pm$ 0.012 0.429 $\pm$ 0.004 0.329 $\pm$ 0.007 0.132 $\pm$ 0.015 0.312 $\pm$ 0.006 0.203 $\pm$ 0.009 0.047 $\pm$ 0.014
UNI [15] 0.369 $\pm$ 0.004 0.282 $\pm$ 0.009 0.116 $\pm$ 0.014 0.438 $\pm$ 0.007 0.331 $\pm$ 0.005 0.127 $\pm$ 0.014 0.288 $\pm$ 0.004 0.183 $\pm$ 0.005 0.054 $\pm$ 0.009
ST-Net [24] 0.372 $\pm$ 0.007 0.284 $\pm$ 0.006 0.108 $\pm$ 0.011 0.426 $\pm$ 0.006 0.327 $\pm$ 0.012 0.125 $\pm$ 0.012 0.309 $\pm$ 0.007 0.198 $\pm$ 0.011 0.057 $\pm$ 0.010
HisToGene [56] 0.366 $\pm$ 0.009 0.279 $\pm$ 0.011 0.134 $\pm$ 0.014 0.422 $\pm$ 0.005 0.318 $\pm$ 0.016 0.146 $\pm$ 0.015 0.306 $\pm$ 0.012 0.194 $\pm$ 0.010 0.064 $\pm$ 0.011
HEST-FT [32] 0.362 $\pm$ 0.010 0.277 $\pm$ 0.011 0.142 $\pm$ 0.011 0.417 $\pm$ 0.006 0.313 $\pm$ 0.013 0.157 $\pm$ 0.019 0.301 $\pm$ 0.008 0.188 $\pm$ 0.007 0.077 $\pm$ 0.007
BLEEP [77] 0.363 $\pm$ 0.008 0.276 $\pm$ 0.014 0.138 $\pm$ 0.013 0.415 $\pm$ 0.007 0.311 $\pm$ 0.015 0.160 $\pm$ 0.009 0.297 $\pm$ 0.009 0.187 $\pm$ 0.009 0.084 $\pm$ 0.005
RankByGene 0.354 $\pm$ 0.007 0.272 $\pm$ 0.009 0.187 $\pm$ 0.010 0.407 $\pm$ 0.006 0.306 $\pm$ 0.012 0.172 $\pm$ 0.008 0.276 $\pm$ 0.008 0.176 $\pm$ 0.005 0.103 $\pm$ 0.004
Table 1: Gene expression prediction results: (a) top 250 genes and (b) top 250 cancer-specific genes.